# The geometry of Deroin–Tholozan representations

### v2

@@ 6.3. Building a hyperbolic cone sphere with prescribed holonomy
We apply Proposition 4 to find a geometric presentation of $\pi_{1}\Sigma^{\prime}$ such that all the triangles in the triangle chain associated to $\rho$ for the standard pants decomposition $\mathcal{B}$ have positive area and don’t overlap with adjacent triangles. The pants curves of $\mathcal{B}$ are represented by the fundamental group elements $b_{1},\ldots,b_{n-3}$. Let $\beta_{1},\ldots,\beta_{n-3}$ and $\gamma_{1},\ldots,\gamma_{n-3}$ be the action-angle coordinates of $\rho$ obtained from the pants decomposition $\mathcal{B}$. Since in the triangle chain associated to $\rho$ no consecutive triangles overlap, it holds that $\gamma_{i}\in[0,\pi]$ for every $i=1,\ldots,n-3$. We can even adapt the geometric presentation to make sure that $\gamma_{i}\in[0,\pi)$ for every $i=1,\ldots,n-3$ because of Remark 6.
@@
Next, from each shared vertex $B_{i}$ of the triangle chain, we draw slits on the northern hemispheres of the adjacent samosas. We choose the slit angles $\phi_{i}\in(0,\beta_{i}/2]$ and $\phi^{\prime}_{i}\in[0,\beta^{\prime}_{i}/2)$ so that $\gamma_{i}=\phi^{\prime}_{i}+\beta_{i}/2-\phi_{i}$. We can always choose the slit angles this way, because the range of possible values of the expression $\phi^{\prime}_{i}+\beta_{i}/2-\phi_{i}$, when $\phi_{i}\in(0,\beta_{i}/2]$ and $\phi^{\prime}_{i}\in[0,\beta^{\prime}_{i}/2)$, is the interval $[0,\pi)$, which coincides with the range of possible values for $\gamma_{i}$. If $\gamma_{i}$ is 00, there’s only one choice: $\phi_{i}=\beta_{i}/2$ and $\phi_{i}^{\prime}=0$. On the other hand, if $\gamma_{i}\in(0,\pi)$, there’s a continuum of possible choices for $\phi_{i}$ and $\phi^{\prime}_{i}$. We give the slits on both sides of $B_{i}$ the same length $\ell_{i}$, so they can be cut and glued together. This turns our sequence of samosas into a samosa assembly whose underlying pants assembly is $\Upsilon$. Finally, because we made sure that $\phi_{i}$ is never equal to 00 and $\phi_{i}^{\prime}$ is never equal to $\beta_{i}^{\prime}/2$, we can pick $\ell_{1},\ldots,\ell_{n-3}$ small enough to ensure that our samosa assembly is a hamantash assembly in $\mathcal{H}_{\alpha,\Upsilon}^{\textnormal{north}}$ (Lemmas 1 and 4). Label the resulting hamantash assembly $H_{\rho}$.
@@ 7. Allowing degenerate configurations
In order to produce more hyperbolic cone structures, such as those in $\overline{\operatorname{Hyp}}_{\alpha}(S)$ (Section 3.2.2), where whole singularities are allowed to coalesce with other singularities, one could study realization of degenerate samosa assemblies (Definition 7). Define $\overline{\mathcal{S}}_{\alpha,\Upsilon}$ to be the space of all samosa assemblies (not only non-degenerate), where slits are now allowed to end at the end of other slits or at fractional cone points. Intuitively, the realization of such samosa assemblies should produce hyperbolic cone structures in $\overline{\operatorname{Hyp}}_{\alpha}(S_{\Upsilon})$. One must nevertheless observe that two slits that are meant to be glued cannot both end on a fractional singularity, because fractional singularities are not allowed to merge in $\overline{\operatorname{Hyp}}_{\alpha}(S_{\Upsilon})$.
@@
Does the realization map from Section 4.3.4 extend meaningfully to a realization map $\overline{R}\colon\overline{\mathcal{S}}_{\alpha,\Upsilon}\to\overline{\operatorname{Hyp}}_{\alpha}(S_{\Upsilon})$? If so, can it be used to chart $\overline{\operatorname{Hyp}}_{\alpha}(S_{\Upsilon})$? Note that even if an extended realization process could produce new hyperbolic cone structures, it wouldn’t produce any new holonomy representations, because every DT representation is already realized by a non-degenerate samosa assembly (Corollary 3).
Interestingly, any degenerate samosa assembly can be seen as a limit of non-degenerate ones by shortening the lengths of the slits that end at fractional singularities and at the end of other slits. One could thus start studying Question 3 by verifying that the realized sequence of hyperbolic cone structures has a limit in $\overline{\operatorname{Hyp}}_{\alpha}(S_{\Upsilon})$. This limit can then be interpreted as the realization of the initial degenerate samosa assembly. Note that all these hyperbolic cone spheres have the same holonomy since we only change slit lengths, highlighting a relation between Questions 2 and 3.
Appendix A Trigonometric formulae
The following classical formulae will be useful for our study of hyperbolic cone metrics. We assume that we’re working with a non-degenerate hyperbolic triangle of side lengths $(a,b,c)$ and interior angles $(\alpha,\beta,\gamma)$.
@@ B.1. Convex combinations of points in the hyperbolic plane
$x\sinh(d(X,Z))=y\sinh(d(Y,Z))$
and the real number $z=x\cosh(d(X,Z))+y\cosh(d(Y,Z))$. The center of mass can be seen as the result of a binary operation: $(Z,z)=(X,x)\ast(Y,y)$. In the special case where the points are the same and only the masses differ, the $\ast$ operation simplifies to $(X,x)\ast(X,x^{\prime})=(X,x+x^{\prime})$.
@@
The mass $c$ can be expressed as $c=\sum_{i}x_{i}\cosh(d(X_{i},C))$ [Sta07, Theorem 3.7].
@@ B.2. Parametrizing hyperbolic segments
The function $t\mapsto Z_{t}$ is a smooth function $(0,1)\to\operatorname{\mathbb{H}}$ whose derivatives all have right and left limits as $t$ approaches 00 and $1$ respectively. As a family parameterized by $Y$, the maps $t\mapsto Z_{t}$ depend continuously on the variable $\eta$ in the $\mathcal{C}^{\infty}$ topology on $\mathcal{C}^{\infty}((0,1),\operatorname{\mathbb{H}})$.
@@ B.3. Parametrizing hyperbolic triangles
$(Z,z)=(X_{1},t_{1})\ast(X_{2},t_{2})\ast(X_{3},t_{3}).$
(We don’t record the value of the mass $z$.) The image of $f$ is the interior of the triangle $(X_{1},X_{2},X_{3})$.

### v3

@@ 6.3. Building a hyperbolic cone sphere with prescribed holonomy
We apply Proposition 4 to find a geometric presentation of $\pi_{1}\Sigma^{\prime}$ such that all the triangles in the triangle chain associated with $\rho$ for the standard pants decomposition $\mathcal{B}$ have positive area and don’t overlap with adjacent triangles. The pants curves of $\mathcal{B}$ are represented by the fundamental group elements $b_{1},\ldots,b_{n-3}$. Let $\beta_{1},\ldots,\beta_{n-3}$ and $\gamma_{1},\ldots,\gamma_{n-3}$ be the action-angle coordinates of $\rho$ for the pants decomposition $\mathcal{B}$. In the triangle chain associated with $\rho$, no consecutive triangles overlap, so $\gamma_{i}\in[0,\pi]$ for every $i=1,\ldots,n-3$. We can even adapt the geometric presentation to make sure that $\gamma_{i}\in[0,\pi)$ for every $i=1,\ldots,n-3$, as discussed in Remark 6.
@@
Next, from each shared vertex $B_{i}$ of the triangle chain, we draw slits on the northern hemispheres of the adjacent samosas. We choose the slit angles $\phi_{i}\in(0,\beta_{i}/2]$ and $\phi^{\prime}_{i}\in[0,\beta^{\prime}_{i}/2)$ so that $\gamma_{i}=\phi^{\prime}_{i}+\beta_{i}/2-\phi_{i}$. We can always choose the slit angles this way, because the range of the expression $\phi^{\prime}_{i}+\beta_{i}/2-\phi_{i}$ with $\phi_{i}\in(0,\beta_{i}/2]$ and $\phi^{\prime}_{i}\in[0,\beta^{\prime}_{i}/2)$ is the interval $[0,\pi)$, which matches the range of $\gamma_{i}$. If $\gamma_{i}$ is 00, there’s only one choice: $\phi_{i}=\beta_{i}/2$ and $\phi_{i}^{\prime}=0$. On the other hand, if $\gamma_{i}\in(0,\pi)$, there’s a continuum of possible choices for $\phi_{i}$ and $\phi^{\prime}_{i}$. We give the slits on both sides of $B_{i}$ the same length $\ell_{i}$, so they can be cut and glued together. This turns our sequence of samosas into a samosa assembly whose underlying pants assembly is $\Upsilon$. Finally, because we made sure that $\phi_{i}$ is never equal to 00 and $\phi_{i}^{\prime}$ is never equal to $\beta_{i}^{\prime}/2$, we can pick $\ell_{1},\ldots,\ell_{n-3}$ small enough to ensure that our samosa assembly is a hamantash assembly in $\mathcal{H}_{\alpha,\Upsilon}^{\textnormal{north}}$ (Lemmas 1 and 4). Label the resulting hamantash assembly $H_{\rho}$.
@@ 7. Allowing degenerate configurations
To parameterize hyperbolic cone structures in the larger space $\overline{\operatorname{Hyp}}_{\alpha}(S)$, where whole singularities are allowed to coalesce with other singularities (Section 3.2.2), one might try to extend the realization process to degenerate samosa assemblies, whose slits are allowed to end at the ends of other slits and at fractional singularities (Definition 7). Intuitively, most degenerate samosa assemblies should have realizations in $\overline{\operatorname{Hyp}}_{\alpha}(S_{\Upsilon})$, but some degenerate samosa assemblies might need to be excluded from the domain $\overline{\mathcal{S}}_{\alpha,\Upsilon}$ of the extended realization map. For example, since fractional singularities aren’t allowed to merge in $\overline{\operatorname{Hyp}}_{\alpha}(S_{\Upsilon})$, it might not be possible to realize samosa assemblies where two slits that end on fractional singularities are meant to be glued together.
@@
Does the realization map from Section 4.3.4 extend meaningfully to a realization map $\overline{R}\colon\overline{\mathcal{S}}_{\alpha,\Upsilon}\to\overline{\operatorname{Hyp}}_{\alpha}(S_{\Upsilon})$, where the domain is expanded to include some degenerate samosa assemblies? If so, can $\overline{R}$ be used to chart $\overline{\operatorname{Hyp}}_{\alpha}(S_{\Upsilon})$?
Remark 7.
Even if an extended realization map could produce new hyperbolic cone structures, it wouldn’t produce any new holonomy representations, because every DT representation is already realized by a non-degenerate samosa assembly (Corollary 3).
Any degenerate samosa assembly can be seen as a limit of non-degenerate ones by shortening the slits that end at fractional singularities and at the ends of other slits. One could thus start studying Question 3 by verifying that the realized sequence of hyperbolic cone structures has a limit in $\overline{\operatorname{Hyp}}_{\alpha}(S_{\Upsilon})$. This limit might then be interpreted as the realization of the initial degenerate samosa assembly. Note that all the realized hyperbolic cone spheres in the sequence would have the same holonomy, since the samosa assemblies in the sequence differ only in their slit lengths. This suggests a relationship between Questions 2 and 3.
Appendix A Trigonometric formulas
The following classical formulas will be useful for our study of hyperbolic cone metrics. We assume that we’re working with a non-degenerate hyperbolic triangle of side lengths $(a,b,c)$ and interior angles $(\alpha,\beta,\gamma)$.
@@ B.1. Convex combinations of points in the hyperbolic plane
$x\sinh\big{(}d(X,Z)\big{)}=y\sinh\big{(}d(Y,Z)\big{)}$
and the real number $z=x\cosh\big{(}d(X,Z)\big{)}+y\cosh\big{(}d(Y,Z)\big{)}$. The center of mass can be seen as the result of a binary operation: $(Z,z)=(X,x)\ast(Y,y)$. In the special case where the points are the same and only the masses differ, the $\ast$ operation simplifies to $(X,x)\ast(X,x^{\prime})=(X,x+x^{\prime})$.
@@
The mass $c$ can be expressed as $c=\sum_{i}x_{i}\cosh\big{(}d(X_{i},C)\big{)}$ [Sta07, Theorem 3.7].
@@ B.2. Parametrizing hyperbolic segments
The function $t\mapsto Z_{t}$ is a smooth function $(0,1)\to\operatorname{\mathbb{H}}$ whose derivatives all have right and left limits as $t$ approaches 00 and $1$ respectively. As a family parameterized by $Y$, the maps $t\mapsto Z_{t}$ depend continuously on the variable $\eta$ in the $\mathcal{C}^{\infty}$ topology on $\mathcal{C}^{\infty}\big{(}(0,1),\operatorname{\mathbb{H}}\big{)}$.
@@ B.3. Parametrizing hyperbolic triangles
$(Z,z)=(X_{1},t_{1})\ast(X_{2},t_{2})\ast(X_{3},t_{3}),$
discarding the mass $z$. The image of $f$ is the interior of the triangle $(X_{1},X_{2},X_{3})$.